Brian 30 years older than his son Travis the total of their ages is 60 find their ages
Found 2 solutions by thesvw, macston:
Answer by thesvw(77)

Let B be Brians age and T be Travis's
B = 30+ T and B+ T=60
Solve them and you wil get Brians age is 60 and Travis 30.

Answer by macston(5194)

.
T=Travis's age; B=Brian's age=T+30yrs
T+B=60 years Substitute for B.
T+T+30 yrs=60 yrs
2T=30 yrs
T=15 yrs ANSWER 1: Travis is 15 years old.
B=T+30yrs=15yrs+30yrs=45yrs ANSWER 2: Brian is 45 years old.
.
CHECK:
T+B=60 yrs
15 yrs+45 yrs=60 yrs
60 yrs=60 yrs